OpenStudy (anonymous):

You had a credit card balance of $358 before you used the card to purchase books for $162, CDs for $81, and a pair of shoes for $92. You then make a payment to the credit card company of $304. Find your updated balance after the payment is processed.

OpenStudy (texaschic101):

ok...you started out with $ 358. now add up all the stuff you bought 162 + 81 + 92 = 335 358 + 335 = 693 -- balance before payment 693 - 304 = 389 -- balance after payment (updated balance)

OpenStudy (anonymous):

oh now i understand thanks :)

OpenStudy (texaschic101):

its a little confusing...but keep in mind, the $358 balance you started with is what you already owe...then you bought more and just added to it. Then , of course, your subtract when you make your payment.

OpenStudy (anonymous):

ya I just got confused now I totally understand thanks to you